The Doppler shift in the wavelength of the sodium line is (5890 angstrom) in the light observed from a distant star is 100 angstrom. To determine the speed at which the star is receding.
1
Expert's answer
2019-12-26T11:59:04-0500

The change in the received frequency due to the Doppler shift can be written as follows:

"\omega = \frac{\omega_0}{1 + \frac{v}{c}}" ,

where we take into account the fact that the star is receding from the observer. Hence,

"v=c\left(\frac{\omega_0}{\omega}-1\right)" .

Utilizing the relationship between the angular velocity and wavelength "\lambda = \frac{2 \pi c}{\omega}" , we derive:

"v=c\left( \frac{\lambda}{\lambda_0}-1\right)=c \left(\frac{\lambda_0 + \Delta \lambda}{\lambda_0}-1\right) = c \frac{\Delta \lambda}{\lambda_0}"

Substituting the numerical values, we obtain (answer is measured in the number of speed of light):

"v = c \frac{100}{5890} \approx 0.017c"

or substituting the speed of light:

"v \approx 5.1 \times 10^6" m/s